Fermi National Laboratory
 

Operation Summary for June 29, 2011 - July 1, 2011

 

Wednesday June 29

The day shift began with Booster down for RF work, with Operations monitoring the Tevatron (TEV) store 8848 (110.05E30), the Recycler stash (332.1E10), and the Antiproton Source (Pbar) in access waiting for work, with the Main Injector (MI) holding on to the Recycler stash, with Linac conducting safety system tests, I- Source maintenance, and working on RF stations, with the NuMI and MiniBooNE experiments off due to Booster being down, but conducting maintenance work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Interlock technicians reported at 8:47 AM that they had completed the Linac safety system testing. Linac personnel reported at 9:48 AM that they had completed work on LRF5 and were beginning work on LRF4. Preaccelerator personnel reported at 10:05 AM that they had completed work on the I- Source. LRF4 work finished at 10:31 AM and Linac was ready for beam by 11:38 AM. Booster was ready for beam by 12:41 PM. Operations began sending beam to NuMI at 12:49 PM. Pbar personnel reported at 12:53 PM that they were having problems with a quadrupole magnet power supply (Q404); an EE Support person investigated. Operators reported at 1:19 PM that the MiniBooNE horn power supply would not turn on; operators asked EE Support personnel to investigate. Pbar resumed stacking at 1:40 PM. Operations terminated store 8848 at 2:34 PM and then prepared D0 for access. Operations established beam to MiniBooNE at 3:32 PM.

During the evening shift, D0 came out of access at 4:54 PM. Operators put the TEV through a dry squeeze at 5:30 PM and through a wet squeeze at 6:10 PM. Operations put the TEV into shot setup at 6:34 PM and established store 8851 at 7:35 PM with an average initial luminosity (AIL) of 398.42E30. Operators tuned the 750 KeV line at 10:58 PM. Operators reported at 11:46 PM that a Booster vacuum pump (IPRF17) tripped off and took down BRF7 & 8; the stations reset, but BRF8 began tripping off and sparking. Operators bypassed BRF8 and brought BRF19 on in its place.

Thursday June 30

The midnight shift began with Operations monitoring store 8851 (169.12E30) and the stash (153.4E10), with Pbar stacking (21.0E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators reported, at 12:33 AM, that the expansion tank for the NuMI absorber LCW system had a level drop; not enough to cause an alarm, but operators notified NuMI personnel anyway. They will investigate in the morning.

During the day shift, Operations terminated store 8851 at 10:06 AM and put the TEV into shot setup. At 11:19 AM, the TEV low beta magnets at TEV sectors BB and DA quenched after the last injection of antiprotons. Operators managed to dump the ramp before the cryo magnets at A4 and C4 could quench. Operators reported that the cause was due to a TEV kicker at E48 that misfired. TEV personnel reported that the initial cause of the quench was due to the spark gap air compressor
tripping off and lowering the spark gap pressure. The Run Coordinator determined at 11:57 AM that this would be a good time to continue with TEV safety system testing. These tests included F-Sector, which halted beam to MI for a short time.

During the evening shift, Operations reported at 4:03 PM that a new Meson FTBF experiment T-1015 had received permission to take beam. Operations established beam to T-1015 around 5:30 PM and disabled it around 6:30 PM. Operators put the TEV through a dry squeeze at 5:43 PM and through a wet squeeze at 7:02 PM. Operations put the TEV into shot setup around 10:35 PM.

Friday July 1

The midnight shift began with the TEV in shot setup, with Operations monitoring the stash (4.7E10), with Pbar stacking (48.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established store 8854 at 12:17 AM with and AIL of 157.89E30. Operations established beam to T-1015 at 4:15 AM. Operators reported short downtimes due to LRF5.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/01/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory